# Predicting wide binaries and deviations from standard gravity using machine learning algorithms

### v2

@@ II.3 Clustering and Nearest Neighbour Search
Given a dataset $\mathcal{D}=\{x_{1},x_{2},\ldots,x_{n}\}\subset\mathbb{R}^{d}$ and a query point $q\in\mathbb{R}^{d}$, the goal of nearest neighbour search (NNS) is to find the point $x^{*}\in\mathcal{D}$ minimizing the distance to $q$, i.e., $x^{*}=\arg\min_{x\in\mathcal{D}}\,\text{dist}(q,x)$. Common choices for the distance function include Euclidean distance $\sqrt{\sum_{i=1}^{d}(q_{i}-x_{i})^{2}}$, cosine distance $1-\frac{q\cdot x}{\|q\|\,\|x\|}$, or other task-specific metrics [27]. While naive search requires $\mathcal{O}(n)$ comparisons and becomes expensive for large datasets, efficiency can be improved using methods like KD-Trees (in low dimensions), approximate nearest neighbour (ANN) techniques such as FAISS or HNSW (suitable for high-dimensional data), or locality-sensitive hashing (LSH) to reduce search complexity.
@@
$D_{\mathrm{3D}}=\sqrt{(x_{A}-x_{B})^{2}+(y_{A}-y_{B})^{2}+(z_{A}-z_{B})^{2}}.$
For each binary system, we search for its nearest neighbouring binary system in this 3D space. The resulting nearest neighbour distances provide a quantitative measure of local stellar density and can be used to flag potentially non-isolated binaries. The distribution of nearest neighbour distances also informs selection criteria: binaries with anomalously small separations to other systems may be excluded or treated with caution, as they are more likely to be influenced by dynamical interactions or unaccounted-for hierarchical structure. In this study, the NNS results are further used in conjunction with clustering algorithms (e.g., DBSCAN) to confirm group memberships and validate the statistical independence of selected binary systems.
III Wide binary classification problem
III.1 The wide binary catalogue
In a wide binary, both the components have the same age and composition which make them fit for astronomical analysis [8]. Moreover, due to their large separations, they help to understand the stellar disk formation in low-density areas [36]. While a wide binary is viewed as two point sources in the sky, chance alignment poses a challenge to the accurate identification of an authentic binary system. For years, the problem of increase in number of chance alignments with the increasing separation has been dealt with through different approaches like the inclusion of proper motion [37, 38] and using parallaxes and radial velocities [39]. The subsequent Gaia data releases [40, 41, 42] has revolutionized the construction of wide binary catalogues by dramatically increasing sample sizes and enabling more precise measurements of parallaxes and proper motions. Based on Gaia DR2 data, several catalogues were formed with different cuts on the separation and parallaxes [43, 44, 45]. Similarly, from the Gaia eDR3 dataset, wide binary catalogues have been generated in [8, 20]. The objective of our work is to predict wide binary pairs from raw Gaia DR3 dataset using a machine learning model trained on the existing catalogue in [8, 46].
@@ III.1 The wide binary catalogue
  1. 1.
    Angular separation condition
    $\theta(/arcsec)\leq 206.265\times\bar{\omega}(/mas)$ (1)
    where $\bar{\omega}$ is the parallax of the star with the brighter G magnitude and $\theta$ is the angle subtended by the arc connecting the two stars. The upper limit on the separation is called the Jacobi radius beyond which the Galactic tidal field becomes comparable to the gravitational attraction of the two stars and is given as $r_{J}\thickapprox 1.35pc\times(\frac{M_{tot}}{M_{o}})$ (Binney et.al [47]).
  2. 2.
    Parallaxes consistent within 6 sigma:
[…]
    $|\bar{\omega}_{1}-\bar{\omega}_{2}|<6\sqrt{\sigma_{\bar{\omega},1}^{2}+\sigma_{\bar{\omega},2}^{2}}$ (2)
@@
  3. 3.
    Orbital proper motion:
    Besides parallax, proper motion, i.e, the motion observed from Earth, also plays a vital role in determining the behaviour of a star/a system. The difference in the proper motion of the two stars are required to be consistent with a bound Keplarian orbit. This suggests that for a system of total mass $5M_{o}$, the projected velocity difference would be
    $\Delta\mu\leq\Delta{\mu}_{orbit}+2\sigma_{\Delta\mu}$ (3)
    where
    $\begin{split}\Delta\mu={[(\mu_{\alpha,1}^{*}-\mu_{\alpha,2}^{*})+(\mu_{\delta,1}-\mu_{\delta,2})^{2}]^{1/2}}\\ \Delta\mu_{orbit}(mas/yr)\leq 0.44(\bar{\omega}(mas))^{3/2}(\theta(arcsec))^{-1/2}\\ \sigma_{\Delta\mu}=\frac{1}{\Delta\mu}{[(\sigma_{\mu_{\alpha,1}}^{*}-\sigma_{\mu_{\alpha,2}}^{*})\Delta\mu_{\alpha}^{2}+(\sigma_{\mu_{\delta,1}}-\sigma_{\mu_{\delta,2}})\Delta\mu_{\delta}^{2}]}^{1/2}\end{split}$ (4)
Subsequent steps include dissolving clusters and cleaning the background to get unbound systems out of moving groups and star clusters, to finally generate the catalogue of labelled wide binaries.
III.2 Methodology
@@
The catalogue chosen for marking the WBS is a subset of the El Badry catalogue. The code for generating the catalogue can be found in the following Zenodo repository: Wide-binaries-from-Gaia-eDR3 [46] written for the paper ”A Million WBS from GAIA eDR3” by El Badry et al. [8].
@@
The first step is the process of data extraction from the GAIA data archive. The data is then filtered and made into a catalogue of Wide Binary Systems (WBS) using Astronomical Data Query Language (ADQL). This task is achieved by systematically following the guidelines mentioned in the work done by Banik et al. [48]. The next step involves marking the obtained WBCs in the raw dataset. This is done so as to create a label that acts as the target variable of the ML model, and the other features in the dataset are the predictors that are used to train the ML model. The next step is the pre-processing and filtering of data to choose the optimal features for ML classification. The positional information, such as right ascension and declination, is intentionally taken out so as to avoid overfitting. After pre-processing, the task of implementing the ML model and performing accuracy measurements is carried out. According to the accuracy requirements, the hyperparameter values are tuned and the most optimal set is selected. Finally, the ML model is used the predict all the WBS in the dataset and then the WBS are paired together using clustering algorithms.
@@
    Additional steps that can be performed at this stage include: PCA reduction, correlation-based filtering: only including the most correlated features or setting some correlation cutoff.
@@
III.3 Performance Report and Results
@@
The confusion matrices also showcase the low detection rate of the base models, which is improved by the SMOTE-balanced models as observed in the FIG. 3(a) and FIG. 3(b). A detailed plot for all the tested algorithms is given in the appendix A.
@@
IV Deviations from standard gravity as an anomaly detection problem
IV.1 Deviations from standard gravity
In context of machine learning, we treat the deviations from standard gravity as an anomaly detection problem for which we require a dataset to train the ML model. We use the catalogue generated by Chae as our training/testing/validation dataset [20]. This catalogue has the binary pairs that are within $80pc$ distance and $200pc$ distance separately from the point of observation. However, for the larger distance, measurement uncertainties increase, while for the smaller distance, the dataset is smaller in number. Moreover, all the binary pairs are of ”MSMS” class. An ”MS” type star can be a giant, a subgiant, a pre-main sequence star or a brown dwarf. Some other constraints of this catalogue are:
  • $|d_{A}-d_{B}|<3\sqrt{\sigma_{d_{A}}^{2}+\sigma_{d_{B}}^{2}}$
  • Relative errors of PM $<0.01$
  • Clean range $144<M_{G}<$ 14 and strict range $124<M_{G}<$ 12
[…]
$log_{10}(\frac{M_{*}}{M_{o}})=\sum_{i=0}^{10}a_{i}(M_{G})^{i},$ (5)
where the coefficients $a_{i}$ corresponding to both $V$ and $J$ bands as obtained in Ref. [20] are given in table 3. The mass is expressed in solar mass units $M_{o}$ and $M_{*}<1.2M_{o}$ for our purposes.
coefficients for $M_{V}$-based $M_{G}$ for $M_{J}$-based $M_{G}$
[…]
Table 3: Coefficients of the polynomial expression for the mass-magnitude relation as given by Chae [20]. The coefficients are applicable for both $M_{V}$ and $M_{J}$ based G-band magnitudes. The mass is expressed in solar mass units $M_{o}$.
IV.2 Labelling deviations
In the training of ML model, we intentionally eliminated the positional information so as to avoid any spatial bias. In literature [18, 9], the deviation from Newtonian gravity has been inferred using the $\tilde{v}$ parameter. For the training of our ML model to identify anomalous systems, we employ this parameter as our primary input. $\tilde{v}$ is given as
$\tilde{v}=\frac{v_{p}}{\sqrt{GM_{tot}/s}}$ (6)
$v_{p}$ is the plane-of-sky velocity given as
$v_{p}=4.7404\times 10^{-3}km/s\times\Delta\mu_{rel}\times d$ (7)
$d$ is the distance in pc to the binary system and $\Delta\mu_{rel}$ is the plane-of-sky proper motion given in section III.1. $M_{tot}$ is the total mass of the system and $s$ is the separation between the two stars.
While the usual practice is to consider the projected components of the velocity and separation, we make use of the 2D plane-of-sky velocity($v_{p}$) and the separation(s) estimated while making the catalogue itself that requires no additional computational steps. We note that $\tilde{v}$ or $v_{p}$ are not accurate measures of deviations. In fact, Chae [20] has defined new parameters like $\delta_{newt-obs}$ which are more accurate but require using MCMC simulations. For our work, however, we use the $\tilde{v}$ parameter defined in Ref. [18] because (i) for training an ML model, the computational cost is optimised and once the model is trained, the anomalies beyond our limits can also be calculated. (ii) our model is trained on the MC realised results for Newtonian gravity [20], thus we are aiming at more accurate results through the model. In summary, the MCMC procedure, being computationally expensive, is discarded. However, its results are used for the training of the ML model, which analyses the $\tilde{v}$ parameters to give anomalies.
The first step is to calculate the magnitude of the plane-of-sky relative proper motion:[20]
$\Delta\mu_{\text{rel}}=\sqrt{(\mu_{\alpha,\mathrm{A}}^{*}-\mu_{\alpha,\mathrm{B}}^{*})^{2}+(\mu_{\delta,\mathrm{A}}-\mu_{\delta,\mathrm{B}})^{2}}.$ (8)
where ($\alpha$) and ($\delta$) are the observed right ascension and the declination components of the proper motions (PMs) of the two components of a binary respectively, $(\mu_{\alpha,\mathrm{A}}^{*},\mu_{\delta,\mathrm{A}})$ and $(\mu_{\alpha,\mathrm{B}}^{*},\mu_{\delta,\mathrm{B}})$.
The magnitude of the plane-of-sky relative velocity is given by: [20]
$v_{p}=4.7404\times 10^{-3}~{}\mathrm{km\,s^{-1}}\times\Delta\mu_{\text{rel}}\times d,$ (9)
where $d$ is the distance in parsecs (pc) to the binary system, and all proper motion values are given in units of milliarcseconds per year (mas yr-1). For $d$, we take the error-weighted mean of the distances to components A and B, denoted $d_{A}$ and $d_{B}$, respectively. The value of the Newtonian circular velocity is given by: [20]
$v_{N}=\sqrt{\dfrac{GM_{\mathrm{tot}}}{s}}$ (10)
where $v_{p}$ is the projected velocity given by Equation (4), $M_{\mathrm{tot}}$ is the total mass of the system, $s$ is the projected separation between the two components, and $G$ is Newton’s gravitational constant. The ratio of two projected quantities $\tilde{v}$ is given by: [20]
$\tilde{v}=\frac{v_{p}}{v_{N}},$ (11)
The $\tilde{v}$ values of the actual data are compared with those of simulated values in a gravity theory (Newtonian gravity). A new parameter called delta ($\Delta_{dev}$) given by:
$\Delta_{dev}=\tilde{v}_{obs}-\tilde{v}_{newtonian}$ (12)
The deviations lying outside ±$\pm 3\sigma$ of $\Delta_{dev}$ are marked as deviations in the actual dataset.
The following scheme was followed for determining deviations as mentioned in the analysis done by Chae et al., after a few modifications. [20]:
Start ProcessLoad Filtered Chae DatasetLoad Chae’s Newtonian Simulation DatasetCalculate $v_{p}$ and $v_{N}$alculate $\tilde{v}_{obs}$alculate $v_{p}$ and $v_{N}$alculate $\tilde{v}_{newtonian}$alculate the $\Delta_{dev}$ ParameterRemove Deviations Inside ±$\pm 3\sigma$Mark Deviations in the Chae DatasetEnd Process
Figure 6: Calculation flowchart
The deviations marked through the above-mentioned process are then used for training ML models.
IV.3 Methodology
We use a supervised ML approach to identify gravitational anomalies present in the data, given a WBS catalogue. The WBS catalogue chosen for training the ML models was a subsection of the El Badry catalogue, generated by varying some of the given parameters to better estimate WBS. [46]. This catalogue was further filtered to only include the ”SOURCE_ID”s of the WBS present in Chae’s analysis. [20]
A supervised ML approach was followed for this task because the dataset is very sparse, and it is proven in the literature [23] that unsupervised learning approaches do not perform satisfactorily and require a labelled dataset approach to guide the model. A crude implementation of the unsupervised approach (Deep Neural Network) was also followed, but it was unable to identify the deviations.
For marking the deviations, the methodology mentioned in section IV.2 (Fig.6) was followed to obtain the deviating ”SOURCE_IDs”. This was achieved by using the data from 2 datasets: Chae’s actual dataset and Chae’s Newtonian simulation dataset. After the determination of the IDs, the ML best models were trained by marking the deviating IDs in the El Badry catalogue.
The best-trained ML models were then exported and used to predict deviating WBS in the entire El Badry catalogue. There is a provision for choosing the optimal model for predicting anomalies, namely, the most accurate model and the highest recall model. The most accurate model picks out the anomalies with the highest accuracy, while the highest recall model predicts almost all the anomalies, although it generates many false positives and is comparatively less accurate.
Load El Badry CatalogueIdentify Deviating IDs (See Fig. 6)Mark Deviating IDs in El Badry CataloguePreprocessing (SMOTE / PCA / Raw / Correlation)Train ModelsK-Fold Cross ValidationSave Best ModelsEnd Process
Figure 7: Methodology for Predicting Deviations
The following was the flow of the program for the problem:
  1. (i)
    Loading: Three datasets have to be loaded - the modified El Badry catalogue, the Chae dataset and the Newtonian simulation Chae dataset. The training of the ML models would be performed on the modified El Badry catalogue. The other two would be used to determine the deviating IDs as mentioned in the section IV.2 (FiG.6).
  2. (ii)
    Identifying deviations: For both the Chae dataset and the Newtonian simulation Chae dataset, $v_{p}$, $v_{N}$, $\tilde{v}$ and $\Delta_{dev}$ are calculated as mentioned in section IV.2. Deviations lying inside ±$\pm 3\sigma\Delta_{dev}$ are removed, and an additional cutoff is placed on the projection separation distance, which should be greater than 7 kAU. After imposing these constraints, the ”SOURCE_ID”s of deviations are exported.
  3. (iii)
    Labelling: The exported IDs of the deviations from the previous step are matched with the El Badry catalogue and the deviations are marked in the El Badry Catalogue.
  4. (iv)
    Preprocessing: A check for NULL value containing columns was performed, and those columns were removed, and a filtered dataset was formed. The models were trained on the datasets that were preprocessed using the following techniques: correlation-based filtering, SMOTE, and PCA reduction. Correlation-based filtering techniques involve filtering out the uncorrelated or less correlated features to the target in the dataset. The two methods used for our model were to only include the most correlated features or to set a correlation cutoff on the features. The SMOTE-balanced dataset was generated from the filtered dataset to be used for training the ML models, along with the filtered dataset. The SMOTE-balancing was done to overcome the issue of class imbalances. The dimensionality of the data in the filtered dataset was reduced from 157 variables to 56, 10 and 13 (3 different PCA-reduced datasets) using PCA reduction on the filtered dataset, top correlation dataset and correlation cutoff dataset. A hybrid reprocessing combination of the above-mentioned techniques was also carried out to generate different datasets. Examples include SMOTE on correlation-filtered datasets and PCA on SMOTE-balanced datasets.
  5. (v)
    Training: A variety of differently preprocessed datasets and a variety of ML algorithms were used in the training phase. The data distributions are highly sparse for this problem as well (refer to subsection IV.1). Therefore, SMOTE was utilised to handle the class imbalances.
    Several machine learning algorithms were trained including: Random Forest Classifier (RFC), Logistic Regression (LR), Support Vector Machine with a Radial Basis Function (RBF) kernel (SVM RBF), Decision Tree Classifier (DTC), K-Nearest Neighbors (KNN), Naive Bayes (NB), and the Bagging Classifier.
  6. (vi)
    K-Fold Cross Validation: The training dataset can be dived into $k$ number of folds ($k=5$ in our analysis). Different hyperparameter values for the Random Forest Classifier and the Balanced Random Forest Classifier with 3 different tuning methods. Firstly, in the accuracy tuning method, the hyperparameters that result in the most accurate model are selected. Secondly, in the F1 score tuning method, the hyperparameters that result in the model that gives the least number of misclassifications are selected. And lastly, in the recall tuning method, the hyperparameters that result in the model that predicts the most number of true positives are selected.
  7. (vii)
    Save Best Models: The best-performing models were saved, and the parameter files were exported. The user has the option to choose from the following two best models, since there is a trade-off between the number of true positives detected and the accuracy of the prediction: the most accurate model or the highest recall model. The most accurate model was a Random Forest Classifier trained on a dataset that was preprocessed to have a correlation cutoff value at 0.05. This means only the features/columns that have a correlation greater than this value were kept in the dataset, and the other features were not used for training. The model achieved an accuracy of 99.4759%. The highest recall model was a Balanced Random Forest Classifier, which was trained by $k$-Fold Cross Validation (5 folds) on a dataset that was preprocessed to have a correlation cutoff value at 0.05. The model had a true positive rate of 98.64%, although there was a reduction in accuracy because of the false positives.
IV.4 Performance Report and Results
Many different models trained on differently preprocessed data were evaluated, and the best trained models are as follows:
  1. (i)
    Most accurate models: The model with the highest accuracy was a Random Forest Classifier (RFC) trained on a dataset that was preprocessed to have a correlation cutoff value at 0.05. The model achieved an accuracy of 99.4759%. Refer to TABLE 4 and TABLE 5 for the performance metrics and the classification analysis, respectively.
  2. (ii)
    Highest recall model: The highest recall model was a Balanced Random Forest (BRF) Classifier, which was trained by $k$-Fold Cross Validation (5 folds) on a dataset that was preprocessed to have a correlation cutoff value at 0.05. The model had a true positive rate of 98.64%, although there was a reduction in accuracy because of the false positives. Refer to TABLE 4 and TABLE 5 for the performance metrics and the classification analysis, respectively.
Table 4: Performance Metrics of the most accurate model and the highest recall model (both trained on the correlation cutoff dataset) for anomaly detection
Algorithms Precision Recall F1 score Accuracy
RFC (most accurate) 0.838710 0.527027 0.647303 0.994759
BRF (highest recall) 0.105797 0.986486 0.191099 0.923788
Table 5: Classification Metrics of the most accurate model and the highest recall model (both trained on the correlation cutoff dataset) for anomaly detection
RFC (most accurate) 78 52.702703 85 57.432432
BRF (highest recall) 146 98.648649 1236 835.135135
The confusion matrices also showcase the low detection rate of the base models, which is improved by the SMOTE-balanced models as observed in the FIG. 8(a) and FIG. 8(b).
(a) Most accurate model
(b) Highest recall model
Figure 8: Confusion matrices for the most accurate model and the highest recall model
There is a trade-off between the accuracy of the model and the number of deviations predicted/detected. Therefore, the choice of the model is left to the user according to their specific needs. The most accurate model predicts anomalies accurately, and hence it gives a list of deviations that are most likely true anomalies, as there are very few false positives. In contrast, the highest recall model predicts all the anomalies present in the dataset but there are a significant number of deviations.
V Conclusions
We have applied machine learning (ML) techniques to the problems of detection of wide binaries and their potential deviations from standard gravity. Wide binary stars—stellar pairs separated by hundreds to thousands of astronomical units—are crucial astrophysical laboratories for testing gravitational theories, including potential deviations from Newtonian dynamics at large separations.
While conventional statistical techniques for detecting wide binaries are computationally expensive, relying on Monte-Carlo simulations and complex probabilistic analyses to rule out chance alignments, machine learning based approach offers a scalable alternative where techniques such as clustering algorithms and nearest neighbour search are used to efficiently predict binaries from noisy background populations. In our implementation, the catalogue generated by El-Badry et al.[46] is the primary training dataset for the analysis of Sec. III. However the trained model can be used to predict pairs of a wide binary system directly from the raw Gaia DR3 data.
The second major focus of this work is detecting the recently reported systematic deviations from standard gravity in the low-acceleration limit of wide binaries [10, 20] by framing it as a supervised anomaly detection problem, aimed at identifying non-standard astrophysical systems, particularly deviations from Newtonian gravity. For this purpose, the catalogue and MCMC simulation data from Ref. [20] have been used as training data. Several supervised and semi-supervised learning models and varied preprocessing techniques like SMOTE, correlation-based filtering, PCA reduction, along with K-Fold Cross Validation were employed to uncover anomalies in the data.
Unlike the results of Ref. [20], however, we use a less accurate parameter $\tilde{v}$ to quantify (and label) the deviations in the training data to avoid the computatinally heavy step of obtaining 3D velocities through Monte-Carlo deprojections. To compensate for this, we use the Newtonian simulation results of Ref. [20] to train the models. However, it turns out the trained models do not rely on computationally complex algorithms like Markov Chain Monte Carlo (MCMC) simulations and can predict deviations very efficiently.
The publicly available tool (hosted at https://github.com/DespCAP/G-ML) developed as part of this work can be used to generate a catalogue of Wide Binary Stars quickly and fairly accurately from the raw GAIA source data. The type of ML model to be used and its hyperparameters, the kind of preprocessing techniques to be used, and clustering criteria are all tunable by the user. Our code also allows for tuning the marking process in the training phase, which enables the user to substitute a better alternative which can more accurately determine the deviations. Therefore, with the invocation of the ML algorithms, even computationally complex parameters ($\tilde{v}$ here) can be used for training and then the trained model(s) can be used to predict without explicitly using the computationally complex parameters. There is also a provision of importing the pretrained model parameters so that the training phase is skipped, and the user can directly use the tool to predict WBS and make catalogues. The anomaly detection tool can be used to roughly determine the deviating systems from standard gravity, given a catalogue. The provision of importing the pretrained model parameters is available here as well. There are two pretrained options available to the user: (i) the most accurate model or (ii) the highest recall model.

### v3

@@ 2.3 Clustering and Nearest Neighbour Search
Given a dataset $\mathcal{D}=\{x_{1},x_{2},\ldots,x_{n}\}\subset\mathbb{R}^{d}$ and a query point $q\in\mathbb{R}^{d}$, the goal of nearest neighbour search (NNS) is to find the point $x^{*}\in\mathcal{D}$ minimizing the distance to $q$, i.e., $x^{*}=\arg\min_{x\in\mathcal{D}}\,\text{dist}(q,x)$. Common choices for the distance function include Euclidean distance $\sqrt{\sum_{i=1}^{d}(q_{i}-x_{i})^{2}}$, cosine distance $1-\frac{q\cdot x}{\|q\|\,\|x\|}$, or other task-specific metrics [Cover and Hart(1967)]. While naive search requires $\mathcal{O}(n)$ comparisons and becomes expensive for large datasets, efficiency can be improved using methods like KD-Trees (in low dimensions), approximate nearest neighbour (ANN) techniques such as FAISS or HNSW (suitable for high-dimensional data), or locality-sensitive hashing (LSH) to reduce search complexity.
@@
$D_{\mathrm{3D}}=\sqrt{(x_{A}-x_{B})^{2}+(y_{A}-y_{B})^{2}+(z_{A}-z_{B})^{2}}.$ (2)
For each binary system, we search for its nearest neighbouring binary system in this 3D space. The resulting nearest neighbour distances provide a quantitative measure of local stellar density and can be used to flag potentially non-isolated binaries. In this study, the NNS results are further used in conjunction with clustering algorithms (e.g., DBSCAN) to confirm group memberships and validate the statistical independence of selected binary systems.
It is important to note that the inferred three-dimensional distances, $D_{3D}$, are subject to uncertainties and should not be interpreted as error-free quantities. In the present work, however, these distances are not employed as precise physical measurements but rather to facilitate the ML training process. As such, the method is not strongly dependent on the exact value of $D_{3D}$ as reported in terms of accuracy values in our results, and moderate deviations within the associated confidence intervals do not significantly affect the training outcome. In principle one could add simulated noise in the dataset, however we leave it for future work.
3 Wide binary classification problem
3.1 The wide binary catalogue
In a wide binary, both the components have the same age and composition which make them fit for astronomical analysis [El-Badry et al.(2021)El-Badry, Rix, and Heintz]. Moreover, due to their large separations, they help to understand the stellar disk formation in low-density areas [Shaya and Olling(2010)]. While a wide binary is viewed as two point sources in the sky, chance alignment poses a challenge to the accurate identification of an authentic binary system. For years, the problem of increase in number of chance alignments with the increasing separation has been dealt with through different approaches like the inclusion of proper motion [Chaname and Gould(2004), Dhital et al.(2010)Dhital, West, Stassun, and Bochanski] and using parallaxes and radial velocities [Andrews et al.(2017)Andrews, Chanamé, and Agüeros]. The subsequent Gaia data releases [Gaia Collaboration et al.(2018)Gaia Collaboration, A. G. A. Brown, and et al., Gaia Collaboration(2021), Gaia Collaboration et al.(2023)Gaia Collaboration, Arenou, F., and et al.] has revolutionized the construction of wide binary catalogues by dramatically increasing sample sizes and enabling more precise measurements of parallaxes and proper motions. Based on Gaia DR2 data, several catalogues were formed with different cuts on the separation and parallaxes [El-Badry and Rix(2018a), Tian et al.(2020)Tian, El-Badry, Rix, and Gould, Hartman and Lépine(2020)]. Similarly, from the Gaia eDR3 dataset, wide binary catalogues have been generated in [El-Badry et al.(2021)El-Badry, Rix, and Heintz, Chae(2023b)]. The objective of our work is to predict wide binary pairs from raw Gaia DR3 dataset using a machine learning model trained on the existing catalogue in [El-Badry et al.(2021)El-Badry, Rix, and Heintz, El-Badry(2021)].
@@ 3.1 The wide binary catalogue
  • Projected separation condition. The projected separation should follow:
    $s\leq 1pc.$
    Equivalently, in terms of the angular separation $\theta$ and the parallax $\tilde{\omega}$:
    $\theta\penalty 10000\ arcsec\leq 206.265\times\tilde{\omega}\penalty 10000\ mas$ (3)
    where $1\penalty 10000\ mas=10^{-3}arcsec$.
  • Parallax condition due to chance alignment. The condition on the difference between the parallaxes is given by,
    $|\tilde{\omega}_{1}-\tilde{\omega}_{2}|<b\sqrt{\sigma^{2}_{\tilde{\omega},1}+\sigma^{2}_{\tilde{\omega},2}}$ (4)
@@
    In Ref. [El-Badry et al.(2021)El-Badry, Rix, and Heintz], the choice $b=6$ for $\theta<4$ arcsec is adopted to minimize instances of chance alignments.
  • Orbital proper motion. Following Refs. [El-Badry and Rix(2018b), El-Badry et al.(2021)El-Badry, Rix, and Heintz] the condition for the difference in proper motion of the two stars to be consistent with a bound Keplerian orbit, amounts to requiring that all candidate binaries have proper motion differences within $3\sigma$ of the maximum velocity difference expected for a system of total mass $5M_{\odot}$ with circular orbits:
    $\Delta\mu=[(\mu_{\alpha,1}^{*}-\mu_{\alpha,2}^{*})^{2}+(\mu_{\delta,1}-\mu_{\delta,2})^{2}]^{1/2}\leq\Delta{\mu}_{orbit}+3\sigma_{\Delta\mu}.$ (5)
    $\Delta{\mu}_{orbit}$ in the above equation is given by,
    $\Delta\mu_{orbit}[mas/yr]\leq 0.44(\bar{\omega}[mas])^{3/2}(\theta[arcsec])^{-1/2};$ (6)
    and,
    $\begin{split}\sigma_{\Delta\mu}=\frac{1}{\Delta\mu}{[(\sigma_{\mu_{\alpha,1}}^{*}-\sigma_{\mu_{\alpha,2}}^{*})\Delta\mu_{\alpha}^{2}+(\sigma_{\mu_{\delta,1}}-\sigma_{\mu_{\delta,2}})\Delta\mu_{\delta}^{2}]}^{1/2},\end{split}$ (7)
    where, $\Delta\mu_{\alpha}^{2}=(\mu_{\alpha,1}^{*}-\mu_{\alpha,2}^{*})^{2}$ and $\Delta\mu_{\delta}^{2}=(\mu_{\delta,1}-\mu_{\delta,2})^{2}$. Subsequent steps include dissolving clusters and cleaning the background to get unbound systems out of moving groups and star clusters, to finally generate the catalogue of labeled wide binaries.
3.2 Methodology
@@
The catalogue chosen for marking the WBS is a subset of the El Badry catalogue. The code for generating the catalogue can be found in the following Zenodo repository: Wide-binaries-from-Gaia-eDR3 [El-Badry(2021)] written for the paper ”A Million WBS from Gaia eDR3” by El Badry et al. [El-Badry et al.(2021)El-Badry, Rix, and Heintz].
@@
The first step is the process of data extraction from the Gaia data archive. The data is then filtered and made into a catalogue of Wide Binary Systems (WBS) using Astronomical Data Query Language (ADQL). This task is achieved by systematically following the guidelines mentioned in the work done by Banik et al. [Banik et al.(2023)Banik, Pittordis, Sutherland, Famaey, Ibata, Mieske, and Zhao]. The next step involves marking the obtained WBCs in the raw dataset. This is done so as to create a label that acts as the target variable of the ML model, and the other features in the dataset are the predictors that are used to train the ML model. The next step is the pre-processing and filtering of data to choose the optimal features for ML classification. The positional information, such as right ascension and declination, is intentionally taken out so as to avoid overfitting. After pre-processing, the task of implementing the ML model and performing accuracy measurements is carried out. According to the accuracy requirements, the hyperparameter values are tuned and the optimal set is selected. Finally, the ML model is used the predict all the WBS in the dataset and then the WBS are paired together using clustering algorithms.
@@
    Additional steps that can be performed at this stage include: PCA reduction, correlation-based filtering: only including the highly correlated features or setting some correlation cutoff.
@@
It is important to note that the sole purpose of the nearest neighbour search is to find pairs of stars constituting wide binaries within the list of predicted wide binary systems. In general, for example in tests of gravity, nearest neighbour search should also include nearest neighbour single stars to avoid contamination.
3.3 Performance Report and Results
@@
Algorithms TP TP rate (%) Misclass- ifications Misclassifica- tion rate(%)
@@
The confusion matrices also showcase the low detection rate of the base models, which is improved by the SMOTE-balanced models as observed in the FIG. 3a and FIG. 3b. A detailed plot for all the tested algorithms is given in the appendix A.
@@
4 Conclusions
We have applied machine learning (ML) techniques to the problem of the detection of wide binaries. Wide binary stars—stellar pairs separated by hundreds to thousands of astronomical units—are crucial astrophysical laboratories for testing gravitational theories, including potential deviations from Newtonian dynamics at large separations.
While conventional statistical techniques for detecting wide binaries are computationally expensive, relying on Monte-Carlo simulations and complex probabilistic analyses to rule out chance alignments, a machine learning based approach offers a scalable alternative where techniques such as clustering algorithms and nearest neighbour search are used to efficiently predict binaries from noisy background populations. In our implementation, the catalogue generated by El-Badry et al.[El-Badry(2021)] is the primary training dataset for the analysis of Sec. 3. However, the trained model can be used to predict pairs of a wide binary system directly from the raw Gaia DR3 data.
The publicly available tool (hosted at https://github.com/DespCAP/G-ML) developed as part of this work can be used to generate a catalogue of Wide Binary Stars quickly and fairly accurately from the raw Gaia source data. The type of ML model to be used and its hyperparameters, the kind of preprocessing techniques to be used, and clustering criteria are all tunable by the user. Our code also allows for tuning the marking process in the training phase, which enables the user to substitute a better alternative which can more accurately determine the wide binary systems. There is also a provision for importing the pretrained model parameters so that the training phase is skipped, and the user can directly use the tool to predict WBS and make catalogues.